# Inverted-circuit zero-noise extrapolation for quantum gate error mitigation

### v1

@@ IV.4 Runs on quantum computing hardware and effects of twirling
Figure 9: Comparison of inverted-circuit ZNE (IC-ZNE), standard ZNE (sZNE) and unmitigated raw data (Raw) for 50 runs of the Grover circuit on the IBM quantum system ibmq_ehningen without (a) and with (b) randomised compiling by Pauli twirling. Like in Fig. 7, the obtained values of $\langle A_{\rm Grover}\rangle$ are shown in a box plot. The corresponding root-mean-square errors (RMSE) indicating the deviations from the exact value $\langle A_{\rm Grover}\rangle_{\rm ideal}=1$ (horizontal dashed line) are displayed in the accompanying plots at the bottom. The most accurate result – with smallest RMSE and a smaller number of statistical outliers (open circles) – is obtained using IC-ZNE with twirling.
@@
The above results, which show the superiority of IC-ZNE over sZNE, but also their limitations in the case of excessive error rates, were obtained on simulators with depolarising noise. On real quantum hardware, however, the errors have a more complex structure: e.g. correlations or crosstalk between neighbouring qubits [34] or coherent errors can occur alongside incoherent ones. The technique of randomised compiling simplifies this structure by transforming the errors occurring in CNOT gates into an incoherent Pauli channel (see Sec. II.3). Therefore, we now compare the results of 50 extrapolation runs obtained on real hardware (textitibmq_ehningen) with or without randomised compiling. The fitting procedure without randomised compiling is the same as described in Sec. IV.2, but with 16 identical instead of different randomly twirled circuits. Moreover, we apply readout error mitigation based on the method M3 [25], both for determining the expectation values $\langle A\rangle$ and the error strengths $\epsilon$.
@@
Figures 9 depict the raw data without (a) and with (b) twirling gates and the mitigated expectation values constructed by linear extrapolation using sZNE or IC-ZNE, respectively. The boxplot shows that the distance between the first and the third quartile of $A_{Grover}$ (i.e., the size of the box) is highest for IC-ZNE in the case without twirling, but, in both cases (with and without twirling), there are fewer or no outliers (open circles) compared to raw data or sZNE. Moreover, the median expectation value obtained by IC-ZNE is closer to the exact value, resulting in a small root-mean-square error (RSME). As discussed in Sec. IV.2, this excellent performance of IC-ZNE, which provides an estimation essentially without bias, is due to the more accurate determination of the error strength.
@@
The mitigated expectation values shown for the sZNE, both, with and without twirling gates display a better accuracy (i.e., smaller RMSE) than the raw data. Twirling results in a smaller number of outliers, but, somewhat surprisingly, in a stronger deviation of the median from the exact value such that, in total, the RMSE turns out to be almost identical with and without twirling.
@@
In contrast, when using IC-ZNE, it is advantageous to utilise twirling, as it significantly improves the mitigation result and accuracy compared to sZNE and raw data. This can also be seen, when examining the root-mean-square error (RMSE) in Fig. 9 (lower part), where it is evident that, similar to the simulation, IC-ZNE outperforms sZNE.
@@
Figure 10 shows the results of 50 runs of the HHL circuit without (a) and with (b) twirling on the backend. As can be seen in Figure 10(a), the RMSE values for ’Raw’, ’sZNE’, and ’IC-ZNE’ are all in the same order of magnitude. Upon inspection of the median values (horizontal lines within the boxes), it is evident that both mitigation methods appear to worsen the outcome. The sZNE method tends to underestimate the correct solution, while IC-ZNE tends to overestimate it.
@@
If twirling is included, see Fig. 10(b), the accuracy of IC-ZNE improves significantly, with perfect agreement between the median value of $\langle A_{\text{HHL}}\rangle$ and the exact result. Consequently, the root-mean-square error is three times lower for IC-ZNE than for sZNE.
The comparison of the shot overhead of the inverted-circuit method (60 000 shots) with raw data (10 000 shots) and sZNE (30 000) shows a significant difference. However, simulations have demonstrated that even with an increase in the number of shots for raw data and sZNE, IC-ZNE outperforms both methods. From the results presented above, it is evident that, in case of standard ZNE, the deviations from the exact value do not predominantly originate from statistical fluctuations, but rather from a systematic shift of the mean value due to the inaccurate assumption of noise scaling. In contrast, IC-ZNE combined with randomised compiling provides estimations with essentially zero bias, unless the noise strength is too high.
@@ V Conclusions and Outlook
This work introduced the method of Inverted-Circuit Zero-Noise Extrapolation (IC-ZNE), which is based on measuring the error strength using inverted circuits. We presented its theoretical foundation, starting from a precise definition of the error strength $\epsilon$, from which we derived a relation between the error strength and the probability of measuring all qubits again in the initial state $|0\rangle$ after applying the circuit and its inverse. This relation is valid under the assumption of predominantly incoherent errors, which can be justified by randomised compiling. Our results on a simulator and quantum device clearly show the advantage of our method: it always delivers the smallest root-mean square deviation from the error-free value and provides estimations with essentially zero bias, unless the noise strength of the original circuit is too large. The reason for the improved performance is the accurate determination of the scaling of the error strength, which deviates from the assumption made in standard ZNE (sZNE) that the error strength is proportional to the number of gates. Our method can be used for arbitrary circuits and does not require any knowledge of the underlying noise or complex error characterisation.
In the future, it may be beneficial to explore alternative methods for scaling circuits, such as adaptive factors [7]. Furthermore, it remains to be explored how our method can be combined with other techniques, such as verifying whether the error strength of the scaled circuits obtained with probabilistic error amplification (PEA) scales as expected. Additionally, the IC-ZNE method can potentially be used to identify runs with high error rates, allowing for the selection of runs that cannot be mitigated.
@@ Acknowledgements.
The authors acknowledge funding from the Ministry of Economic Affairs, Labour and Tourism Baden-Württemberg, under the project Sequoia - End to End in the frame of the Competence Center Quantum Computing Baden- Württemberg.
@@ Appendix A Adjoint noise channel
In the following, we ignore the error of the single-qubit gates. If we apply randomised compiling, the errors of the CNOT gates are given by Pauli channels (i.e. $C\to\Lambda_{\mathcal{P}}{\mathcal{C}}$):
@@
Therefore, ${\mathcal{E}}_{U^{\dagger}}$ and ${\mathcal{E}}_{U^{\dagger}}^{\dagger}$ may in principle differ for Pauli channels where $p_{i}\neq p_{\pi(i)}$ for all $i$. However, we expect the effect on the fidelity to be small, since the latter mainly originates from the Pauli terms proportional to the identity operator $P_{1}$ (no error), which commutes with CNOT.
@@ Appendix B Transpiled circuits
In Figs. 11 and 12, we illustrate the quantum circuits utilised in this paper, specifically tailored to conform to IBM’s quantum device architecture through transpilation. Both circuit diagrams incorporate at least one SWAP gate, necessitated by the constraints of the device architecture, thereby augmenting the total count of CNOT gates. The Grover circuit showcased in Fig. 11 requires only one SWAP gate, resulting in a cumulative total of 10 CNOT gates. In contrast the HHL circuit depicted in Fig. 12 exhibits a total of 18 CNOT gates, of which 9 are attributed solely to the inclusion of SWAP gates.
Appendix C Error rates
@@
Figure 13: Histogram of the error rates for each CNOT gate on IBM’s cloud device textitimbq_ehningen from August 18th, 2023. The dashed green line represents the median error rate of the device. The green bars represent the CNOT gates used for runs with the HHL circuit ([4, 1],[6, 7], [7, 6],[4, 7],[7, 4],[4, 1],[6, 7]), while the hatched bars represent those used for runs with the Grover circuit ([4, 7],[6, 7],[7, 6]). The reported error rates for the CNOT gates used are lower than the median error rate and are below 1%.
@@
The error rates of all CNOT gates of ibmq_ehningen (from August 18th, 2023) as reported by IBM are displayed in Fig. 13. We observe variations in error rates across different CNOT gates, which can significantly impact the success probability of quantum algorithms. The median error rate, indicated by the dashed green line, is 0.705%. The CNOT gates used for runs involving the HHL circuit are represented by the green bars ([4, 1],[6, 7], [7, 6],[4, 7],[7, 4],[4, 1],[6, 7]), while the hatched bars represent those employed for runs with the Grover circuit ([4, 7],[6, 7],[7, 6]).
@@
Note that these error rates only give a partial description of the noise occurring on the real device (e.g., they do not include crosstalk effects [34]). Therefore, the total error strengths measured by us using the method of inverted circuits ($\epsilon_{0}\simeq 0.15$ for the Grover circuit and $\epsilon_{0}\simeq 0.4$ for the HHL circuit, see the green diamonds in Figs. 4 and 5) are larger than expected from the error rates per CNOT gate displayed in Fig. 13.

### v3

@@ IV.4 Runs on quantum computing hardware and effects of twirling
Figure 9: Comparison of inverted-circuit ZNE (IC-ZNE), standard ZNE (sZNE) and unmitigated raw data (Raw) for 50 runs of the Grover circuit on the IBM quantum system ibmq_ehningen without (a) and with (b) randomized compiling by Pauli twirling. Like in Fig. 7, the obtained values of $\langle A_{\rm Grover}\rangle$ are shown in a box plot. The corresponding root-mean-square errors (RMSE) indicating the deviations from the exact value $\langle A_{\rm Grover}\rangle_{\rm ideal}=1$ (horizontal dashed line) are displayed in the accompanying plots below. The most accurate result – with smallest RMSE and a smaller number of statistical outliers (open circles) – is obtained using IC-ZNE with twirling.
@@
The above results, which show the superiority of IC-ZNE over sZNE, but also the limitations in the case of excessive error rates, were obtained on simulators with depolarizing noise. On real quantum hardware, however, the errors have a more complex structure: e.g. correlations or crosstalk between neighboring qubits [42] or coherent errors can occur alongside incoherent ones. The technique of randomized compiling simplifies this structure by transforming the errors occurring in CNOT gates into an incoherent Pauli channel (see Sec. II.3). Therefore, we now compare the results of 50 extrapolation runs obtained on real hardware (ibmq_ehningen) with or without randomized compiling. The fitting procedure without randomized compiling is the same as described in Sec. IV.2, but with 16 identical, instead of different randomly twirled circuits. Moreover, we apply readout error mitigation based on the method M3 [33], both for determining the expectation values $\langle A\rangle$ and the error strengths $\epsilon$.
@@
For the Grover circuit, Fig. 9 depicts the raw data without (a) and with (b) twirling gates and the mitigated expectation values constructed by exponential extrapolation using sZNE or linear extrapolation using IC-ZNE, respectively. The boxplot shows that the distance between the first and the third quartile of $A_{Grover}$ (i.e., the size of the box) is higher for IC-ZNE in the case without twirling, but, in both cases (with and without twirling), there are fewer or no outliers (open circles) compared to raw data or sZNE. Moreover, the median expectation value obtained by IC-ZNE with twirling is closer to the exact value, resulting in a small root-mean-square error (RSME). As discussed in Sec. IV.2, this excellent performance of IC-ZNE, which provides an estimation essentially without bias, is due to the more accurate determination of the error strength, if the errors are not described only by depolarizing noise.
@@
The mitigated expectation values shown for the sZNE, both, with and without twirling gates display better accuracy (i.e., smaller RMSE) than the raw data. Twirling results in a smaller number of outliers, but, somewhat surprisingly, in a stronger deviation of the median from the exact value such that, in total, the RMSE turns out to be almost identical with and without twirling.
@@
In contrast, when using IC-ZNE, it is advantageous to utilize twirling, as it improves the mitigation result and accuracy compared to sZNE and raw data. This can also be seen, when examining the root-mean-square error (RMSE) in Fig. 9 (lower part), where it is evident that IC-ZNE outperforms sZNE.
@@
Fig. 10 shows the results of 50 runs of the HHL circuit without (a) and with (b) twirling on the backend. As can be seen in Fig. 10(a), the RMSE values for “Raw” and “IC-ZNE” are in the same order of magnitude. However, “sZNE” shows a wider range of values, resulting in a higher RMSE value. Upon inspection of the median values (horizontal lines within the boxes), it is evident that both mitigation methods overestimate the expectation value without twirling.
@@
If twirling is included, see Fig. 10(b), the accuracy of IC-ZNE improves significantly, with perfect alignment between the median value of $\langle A_{\text{HHL}}\rangle$ and the exact result. Consequently, the root-mean-square error is three times lower for IC-ZNE than for sZNE.
Finally, we would like to comment on the additional resources required for IC-ZNE. As compared to standard ZNE, the execution time on the quantum hardware is approximately twice as large due to the additional circuits required for the measurement of error strengths. The fact that, due to addition of the inverse, these circuits are also twice as deep has little impact on the execution time, which, in present IBM devices, is mainly limited by the preparation of the initial state and hence approximately proportional to the total number of shots.
The experiments discussed in Secs. IV.3 and IV.4 were performed with 10 000 shots for the unmitigated raw data, 30 000 shots for standard ZNE and 60 000 shots for IC-ZNE, respectively. Alternatively, one could also perform standard ZNE and IC-ZNE with the same resources by either increasing the number of shots per circuit by a factor of two in case of standard ZNE or reducing it by a factor of two in case of IC-ZNE. Although this would sightly affect the variance of statistical fluctuations observed in Figs. 9 and 10, we expect that the superiority of IC-ZNE compared to standard ZNE on the real hardware would remain, since IC-ZNE (in combination with randomized compiling) provides estimations with essentially zero bias by avoiding systematic shifts due to a more accurate determination of noise scaling.
This expectation is confirmed by simulations shown in Appendix C, showing that the results of IC-ZNE do not strongly deteriorate when using half the number of shots.
@@ V Conclusions and Outlook
This work introduced the method of Inverted-Circuit Zero-Noise Extrapolation (IC-ZNE), which is based on measuring the error strength using inverted circuits. We presented its theoretical foundation, starting from a precise definition of the error strength $\epsilon$, from which we derived a relation between the error strength and the probability of measuring all qubits again in the initial state $|0\rangle$ after applying the circuit and its inverse. This relation is valid under the assumption of predominantly incoherent errors, which can be justified by randomized compiling. Our results in numerical simulations and on a quantum device clearly show the advantage of our method: it invariably delivers the smallest root-mean square deviation from the error-free value and provides estimations with essentially zero bias, unless the noise strength of the original circuit is too large. The reason for the improved performance is the accurate determination of the scaling of the error strength, which deviates from the exponential scaling assumption made in standard ZNE (sZNE). Our method can be used for arbitrary circuits and does not require any knowledge of the underlying noise or complex error characterization.
In the future, it may be beneficial to explore alternative methods for scaling circuits, such as adaptive factors [7]. Furthermore, it remains to be explored how our method can be combined with other techniques, such as dual-state purification [21, 22], or verifying whether the error strength of the scaled circuits obtained with probabilistic error amplification (PEA) scales as expected. Additionally, the IC-ZNE method can potentially be used to identify runs with high error rates, allowing for the selection of runs that cannot be mitigated. The data that support the findings of this study are available from the corresponding author, upon reasonable request.
@@ Acknowledgements.
The authors acknowledge funding from the Ministry of Economic Affairs, Labor and Tourism Baden-Württemberg, under the project Sequoia - End to End in the frame of the Competence Center Quantum Computing Baden- Württemberg.
@@ Appendix A Adjoint noise channel
In the following, we ignore the error of the single-qubit gates. If we apply randomized compiling, the errors of the CNOT gates are given by Pauli channels (i.e. $C\to\Lambda_{\mathcal{P}}{\mathcal{C}}$):
@@
Therefore, ${\mathcal{E}}_{U^{\dagger}}$ and ${\mathcal{E}}_{U^{\dagger}}^{\dagger}$ may in principle differ for Pauli channels where $p_{i}\neq p_{\pi(i)}$ for at least one $i$. However, we expect the effect on the fidelity to be small, since the latter mainly originates from the Pauli terms proportional to the identity operator $P_{1}$ (no error), which commutes with CNOT.
@@ Appendix B Transpiled circuits
In Figs. 11 and 12, we illustrate the quantum circuits utilized in this paper, specifically tailored to conform to IBM’s quantum device architecture through transpilation. Both circuit diagrams incorporate at least one SWAP gate, necessitated by the constraints of the device architecture, thereby augmenting the total count of CNOT gates. The Grover circuit showcased in Fig. 11 requires only one SWAP gate, resulting in a cumulative total of 10 CNOT gates. In contrast the HHL circuit depicted in Fig. 12 exhibits a total of 18 CNOT gates, of which 9 are attributed solely to the inclusion of SWAP gates.
Appendix C Simulation of IC-ZNE with reduced number of shots
In Fig. 13, we compare the simulation results of IC-ZNE obtained with 30.000 and 60.000 shots in total. The quality of the result, measured in terms of the root mean-square error (RSME), is almost the same in both cases.
[…]
Appendix D Device properties
Figure 14: Connectivity (black lines) between 27 qubits (black circles) of IBM’s Ehningen device.
@@ Appendix D Device properties
Figure 15: Histogram of the error rates for each CNOT gate on IBM’s cloud device imbq_ehningen from August 18th, 2023. The dashed green line represents the median error rate of the device. The green bars represent the CNOT gates used for runs with the HHL circuit ([4, 1],[6, 7], [7, 6],[4, 7],[7, 4],[4, 1],[6, 7]), while the hatched bars represent those used for runs with the Grover circuit ([4, 7],[6, 7],[7, 6]). The reported error rates for the CNOT gates used are lower than the median error rate and are below 1%.
This section presents the detailed specifications of the data obtained from IBM’s quantum device, as described in the main text. Fig. 14 illustrates the connectivity of the device ibmq_ehningen, which hosts 27 qubits in a heavy hexagonal lattice.
@@
The error rates of all CNOT gates of ibmq_ehningen (from August 18th, 2023) as reported by IBM are displayed in Fig. 15. We observe variations in error rates across different CNOT gates, which can significantly impact the success probability of quantum algorithms. The median error rate, indicated by the dashed green line, is 0.705%. The CNOT gates used for runs involving the HHL circuit are represented by the green bars ([4, 1],[6, 7], [7, 6],[4, 7],[7, 4],[4, 1],[6, 7]), while the hatched bars represent those employed for runs with the Grover circuit ([4, 7],[6, 7],[7, 6]). The precise error rates for each qubit pair is presented in Table 1.
@@
Note that these error rates only give a partial description of the noise occurring on the real device (e.g., they do not include crosstalk effects [42]). Therefore, the total error strengths measured using the method of inverted circuits ($\epsilon_{0}\simeq 0.15$ for the Grover circuit and $\epsilon_{0}\simeq 0.4$ for the HHL circuit, see the green diamonds in Figs. 4 and 5) are typically larger than expected from the error rates per CNOT gate displayed in Fig. 15.
The single-qubit device properties of ibmq_ehningen as reported by IBM are presented in Table 2.
Table 1: Table of the gate error of the available two-qubit gate pairs for ibmq_ehningen (as of August 18th, 2023).
CX Gate Pair Gate error CX Gate Pair Gate error
25_22 0.00565 10_12 0.00849
22_25 0.00565 12_10 0.00849
20_19 0.00602 7_10 0.02103
19_20 0.00602 10_7 0.02103
14_16 0.00847 19_22 0.00993
16_14 0.00847 22_19 0.00993
18_17 0.00896 21_23 0.00961
17_18 0.00896 23_21 0.00961
14_11 0.00959 13_12 0.00629
11_14 0.00959 12_13 0.00629
12_15 0.00726 26_25 0.00684
15_12 0.00726 25_26 0.00684
8_5 0.0101 3_2 0.00358
5_8 0.0101 2_3 0.00358
1_0 0.01138 7_6 0.00595
0_1 0.01138 6_7 0.00595
1_2 0.00681 16_19 0.00676
2_1 0.00681 19_16 0.00676
8_11 0.01272 18_21 0.00607
11_8 0.01272 21_18 0.00607
24_23 0.00885 4_7 0.00542
23_24 0.00885 7_4 0.00542
4_1 0.00655 14_13 0.00667
1_4 0.00655 13_14 0.00667
8_9 0.03875 24_25 0.00839
9_8 0.03875 25_24 0.00839
18_15 0.00544 3_5 0.00588
15_18 0.00544 5_3 0.00588
Table 2: Table of the single-qubit device properties of ibmq_ehningen (as of August 18th, 2023). The table displays the frequencies, T1 and T2, gate errors, and readout errors for the various qubits, as reported by IBM.
Qubit Frequency / GHz T1 /µs T2 /µs RZ error SX error X error Readout error
Q0 4.961 142.59275 95.47855 0 0.00018 0.00018 0.011
Q1 5.18191 181.00386 70.72969 0 0.00021 0.00021 0.0096
Q2 5.12694 95.06903 9.06901 0 0.0003 0.0003 0.009
Q3 5.26815 99.89562 28.579 0 0.00017 0.00017 0.0229
Q4 5.05357 159.58091 70.14323 0 0.00068 0.00068 0.0123
Q5 5.07116 104.48761 8.15402 0 0.00032 0.00032 0.024
Q6 4.89006 143.66941 122.05597 0 0.00035 0.00035 0.014
Q7 4.97776 123.99653 113.28929 0 0.00021 0.00021 0.0076
Q8 5.17419 84.87928 69.68872 0 0.00104 0.00104 0.0245
Q9 4.9925 112.5169 62.87694 0 0.00573 0.00573 0.0175
Q10 4.83511 191.01486 162.14782 0 0.00039 0.00039 0.0092
Q11 5.11944 92.86756 127.55782 0 0.00027 0.00027 0.017
Q12 4.72549 169.03301 238.81786 0 0.00015 0.00015 0.0117
Q13 4.92598 155.88103 225.73924 0 0.00022 0.00022 0.008
Q14 5.17671 94.24501 258.67336 0 0.00033 0.00033 0.0077
Q15 4.89299 93.54811 176.56123 0 0.00019 0.00019 0.0105
Q16 5.02214 194.67907 191.30807 0 0.00028 0.00028 0.0078
Q17 5.13566 146.76432 22.47776 0 0.00052 0.00052 0.0073
Q18 4.99642 130.38322 228.58117 0 0.00024 0.00024 0.0128
Q19 4.7841 172.28179 77.40924 0 0.00032 0.00032 0.0137
Q20 5.04235 204.49692 220.82418 0 0.00053 0.00053 0.025
Q21 4.93974 126.28787 202.81237 0 0.00026 0.00026 0.0076
Q22 4.72513 58.42632 34.32104 0 0.00025 0.00025 0.0131
Q23 4.80479 190.68715 238.40096 0 0.00043 0.00043 0.0084
Q24 5.07449 233.74912 337.44479 0 0.00015 0.00015 0.0075
Q25 4.95019 198.39525 439.17036 0 0.00018 0.00018 0.0076
Q26 5.15132 190.73366 24.68715 0 0.00015 0.00015 0.0079